OpenStudy (anonymous):

Find the height of a square pyramid with volume 102 m3 and dimensions of base are 6.5 m by 6.5 m.

OpenStudy (blank ):

height = 7.24

OpenStudy (blank ):

h=3 (V/lw) =3·(102/(6.5·6.5)) = 7.2426

OpenStudy (anonymous):

thanks